If a cat repeatedly rubs against your cotton slacks on a dry day, the charge transfer between the cat hair and the cotton can leave you with an excess charge of-3.80 μC. (a) How many electrons are transferred between you and the cat? You will gradually discharge via the floor, but if instead of waiting, you immediately reach toward a faucet, a painful spark can suddenly appear as your fingers near the faucet. (b) In that spark, do electrons flow from you to the faucet or vice versa? (c) Just before the spark appears, do you induce positive or negative charge in the faucet? (d) If, instead, the cat reaches a paw toward the faucet, which way do electrons flow in the resulting spark, from faucet to the cat or vice versa? (a) Number Units
Answer #1

(a)
Given that
charge, q = -3.80 μC = -3.80 * 10^-6 C
No. of electrons
n = q/e = (-3.80 * 10^-6 C) / (-1.6 * 10^-19 C) = 2.375 * 10^13 electrons
(b)
Since you have the excess electrons, the electrons jump from you to the faucet. 1
(c)
Since unlike charges attract, the faucet becomes positively charged. 1
(d)
The cat is charged positively and due to excess of electrons the flow of charge is from the faucet to the cat. 1